Evaluate S Ss y2z2 ds, where S is the part of the cone z = vx2 + y2 that lies between the planes z = 1 and 2 = 2.
Answer #1

I- S S is part of z= 5x²+ y2 2=1 & 2=2. Now ds = 51+ (2284 Zyg 2 dady x 2x 2g = 25x²ty? Joertya y Zy 52²ty2 ds 1 + x² + y2 +Now change in polor form Pat x=rloso Yarfino x² + y2 =22 2= 8 & 2= 2 ī ņ & 8= 2 rardo docdy limit to Q fo ds = 52.rardo I = gI = 27 CL- -Cosao 6 I = resha I = 6 I = + [o-sivo e (str.) The 65 fax-sigt ) -(o- fing)] [(at-o) - (0-0) I = 65 12 I = = 65x2